hiya:

Solve the equation. Show each step of your work. 5^2x+7=125^x 1

hiya:

@Hoodmemes

hiya:

@Vocaloid

hiya:

@Elsa213

hiya:

@carmelle

hiya:

@ramen

hiya:

@mxddi3

hiya:

@dude

hiya:

@nevaehhh

nevaehhh:

The end number '1' is that supposed to be added or subtracted?

hiya:

idk

1 attachment
nevaehhh:

I'm sorry, but I'm not completely sure how to solve it without knowing whether it's subtraction or addition.

hiya:

oh do you know anyone who could help me?

hiya:

just forget about the 1 and slove it for me please

CocoCammron:

5^2*x-7-(125)=0 Step by step solution : STEP 1 : Equation at the end of step 1 : ((52 • x) - 7) - 125 = 0 STEP 2 : Equation at the end of step 2 25x - 132 = 0 STEP 3 : Solving a Single Variable Equation: 3.1 Solve : 25x-132 = 0 Add 132 to both sides of the equation : 25x = 132 Divide both sides of the equation by 25: x = 132/25 = 5.280 One solution was found : x = 132/25 = 5.280